OpenStudy (anonymous):

please solve!! −4|8−x|≤−20

OpenStudy (anonymous):

x less than or equal to 3, or x greater than or equal to 13

OpenStudy (anonymous):

First, divide by -4 on both sides. Remember to flip the inequality. \[|8-x| \ge 5\]Since this is the greater than case, we have to write two separate inequalities after removing the absolute value. \[\left(\begin{matrix}8-x \ge 5 \\ 8-x \ge -5\end{matrix}\right)\]